OpenStudy (anonymous):

4/3x+2=-1 solve for x

OpenStudy (anonymous):

first subtract 2 from each side. then divide both sides by 4/3 which is essentially ultiplying each sides by 3/4

OpenStudy (diyadiya):

Oh Wait is it \[\frac{4}{3x}+2=-1 \ or \ \frac{4}{3}x+2=-1\]

OpenStudy (diyadiya):

@khan_sara_14

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

she wants us to keep our answer in fraction form

OpenStudy (anonymous):